Will you please explain the following question in detail for me. I have the answer, but do not understand the process to get to the answer. Thank you.

4.4 Suppose that a data warehouse for Big University consists of the four dimensions student, course, semester, and instructor, and two measures count and avg grade. At the lowest conceptual level (e.g., for a given student, course, semester, and instructor combination), the avg grade measure stores the actual course grade of the student. At higher conceptual levels, avg grade stores the average grade for the given combination.

(c) If each dimension has five levels (including all), such as “student < major <status < university all”, how many cuboids will this cube contain (including the base and apex cuboids).  

The answer is : The cube will contain 54=625 cuboids.. Please explain this to me in detail. Thank you.
Answer #1

Question:---------- (c) If each dimension has five levels (including all), such as “student < major <status < university all”, how many cuboids will this cube contain (including the base and apex cuboids).  

Answer:-------

Li = 5-1 =4

N=4 dimensions

So, this cube will contain (Li+1)^4 = 625 cuboids